OpenStudy (anonymous):

If 3a-9+6 , what is the value of 4a-1?

OpenStudy (lgbasallote):

rewrite

mathslover (mathslover):

check the question again ...

OpenStudy (anonymous):

sorry the + is suppose to be =

mathslover (mathslover):

you mean : 3a-9=6

OpenStudy (anonymous):

yes

mathslover (mathslover):

so after that add 9 on both sides: 3a-9+9=6+9

OpenStudy (anonymous):

becomes 3a-18=15

mathslover (mathslover):

no.no. -9+9 = 0

mathslover (mathslover):

they have diff. signs

OpenStudy (anonymous):

oh yea, i didnt notice, so really is becomes 3a=15?

mathslover (mathslover):

yes ... and then divide both sides by 3

OpenStudy (anonymous):

becomes 5a

OpenStudy (anonymous):

i ment a= 5

mathslover (mathslover):

\[\large{\frac{3a}{3}=\frac{15}{3}}\] \[\large{\frac{\cancel{3}a}{\cancel{3}}=\frac{\cancel{15}^5}{\cancel{3}^1}}\]

mathslover (mathslover):

a= 5 right

mathslover (mathslover):

now put this value in 4a-1 4(5)-1

mathslover (mathslover):

what r u getting?

OpenStudy (anonymous):

20-1=19?

mathslover (mathslover):

very good ..

mathslover (mathslover):

that is ur answer

mathslover (mathslover):

and also: \[\huge{\mathbb{Welcome}\textbf{To}\mathbb{Openstudy}}\]